1. give an example of number x,y such that |x| + |y| = x+y

2. line through (3,4) perpendicular to the line given by y= 4x-2
Answer #1

Example consider x=2, 4-3 then 100 1-2 and 11-3 - 1x+y)= 2+3 =5 and alty = 2+3=5 | 241 +141 = 2+y Given the line = 40-2 so th